OpenStudy (anonymous):

say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. find a if b = 6 and c = 10

OpenStudy (anonymous):

8

OpenStudy (anonymous):

A^2 + B^2 = C^2 A^2 + 6^2 = 10^2 A^2 + 36 = 100 A^2 = 64 A=8